# thanks: Present Address: Physikalisch-Technische Bundesanstalt, 38116 Braunschweig, Germany

### v1

@@ Abstract
Introduction – Optical atomic clocks based on frequency spectroscopy of dipole-forbidden electronic transitions in isolated, trapped atoms are among the most precise instruments developed, capable of measuring time more precisely than the cesium clocks that currently define the second LudlowOpticalClockReview2015 . Accordingly, optical clock frequency ratios are some of the most accurate measurements BeloyBacon2021 ; HausserIndiumMultiIon2025 , and are used as probes for new physics SafronovaNewPhysicsReview2018 , including time variation of fundamental constants SafronovaFundamentalConstantVariation2019 ; SherrillFundamentalConstantVariation2023 ; violations of local position invariance LangePositionInvariance2021 ; constraints on dark matter BeloyBacon2021 ; KennedyDarkMatterClock2020 ; FilzingerClocksDM2023 ; and general relativity at small scales ChouAlClockRelativity2010 ; BothwelStrontiumEnsembleRelativity2022 . Optical clocks based on single trapped ions BrewerAccuracy2019 and neutral atoms in optical lattices AppeliStrontiumAccuracy024 have reached fractional frequency uncertainties below $10^{-18}$; further advances open new possibilities for these investigations. Additionally, as the scientific community moves towards the redefinition of the second, advances in the state of the art for clock accuracy and stability are critical DimarcqSecondRoadmap2024 .
@@
We probe both the $\mathrm{m_{F}}=+5/2$ and $\mathrm{m_{F}}=-5/2$ ${}^{1}\mathrm{S}_{0}\leftrightarrow^{3}\mathrm{P}_{0}$ transitions, and generate a “virtual” first-order magnetic-field insensitive transition from their mean frequency RosenbandAlClockObservation2007 . Additionally, we alternate probing from opposite directions, with the two probe beams counterpropagating through the same single-mode optical fibers; the average of opposite directions is insensitive to possible first-order Doppler shifts due to ion motion. Both probe laser AOMs are switched on during each cycle, with the inactive direction detuned by 100 kHz. This maintains a nearly identical electric- and magnetic-field environment for probes in either direction and either side of the 27Al+ Zeeman structure.
Decoherence of the 27Al+ clock laser ($\lambda_{L}\simeq 267$ nm) limited previous generations to a 150 ms probe time. We extend the laser coherence by transferring the stability of the JILA cryogenic silicon cavity MateiSiCavity2016 to the clock laser. An ultrastable laser locked to this cavity travels over a 3.6 km path-length-stabilized fiber link to NIST. This fiber link via the Boulder Research and Administrative Network has been described previously BeloyBacon2021 ; for this work, it has been modified as shown in Fig. 1A, such that the cavity-stabilized light serves as reference for an Er:Yb frequency comb FortierCombReview2019 ; NardelliErYbComb2023 . We first pre-stabilize the 27Al+ clock laser with a Pound-Drever-Hall lock to a room-temperature ultra-low expansion glass cavity. With a lower bandwidth $\sim$10 kHz, we then modulate the drive frequency of an acousto-optic modulator to stabilize the beatnote between clock laser and frequency comb. With this improved clock laser stabilization, we can operate the clock using a 1 s probe duration without substantial loss of atom-laser coherence.
@@
We measure the 27Al+ clock stability by comparison to the JILA Sr optical lattice clock AeppliJILASrClock2024 , with results shown in Fig. 1B. The optical lattice clock fractional stability is $<1\times 10^{-16}/\sqrt{\tau/\mathrm{s}}$, meaning the comparison stability of $3.5\times 10^{-16}/\sqrt{\tau/\mathrm{s}}$ is dominated by the 27Al+ clock. To our knowledge, this represents the lowest instability of any ion clock reported to date. In the future, this stability could be improved by further extending the probe time using differential spectroscopy KimDiffspec2023 , or by extending the quantum logic clock to multiple spectroscopy ions CuiScalableQLS2022 ; KellerMultIonConf2024 ; PelzerMultiIonCa2024 .
@@
Secular motion – We evaluate the energy in the secular modes of the two-ion crystal using sideband thermometry LeibfriedQuantumDynamics2003 ; WinelandExperimentalIssues1998 . To quantify the statistical uncertainty and test the repeatability of sideband thermometry at the Doppler limit, we repeat the measurements many times over the course of several months. The results are consistent within the measurement uncertainty with the calculated Doppler limit for our nominal laser geometry and secular mode frequencies. Figure 2 shows the record of measurements for one of the radial “stretch” modes, with the others shown in supplemental material SM .
@@
Blackbody radiation – The 27Al+ clock transition has one of the lowest polarizabilities due to blackbody radiation of any existing atomic clock SafronovaBBRclocks2012 . This was recently measured WeiPolarizability2024 with high precision in a 27Al+-40Ca+ clock, using the Stark shift on the 40Ca+ clock transition as a reference for the intensity of a Stark shifting infrared laser and giving a polarizability of $(6.86\pm 0.23)\times 10^{-42}$ $\mathrm{Jm}^{2}/\mathrm{V}^{2}$. We measure the temperature of our apparatus using a set of six thermocouples - three located on the trap and support structure, and three on the surrounding vacuum chamber - constraining the ion’s blackbody environment to be $24.0\pm 3.3$ C, for a Stark shift due to blackbody radiation of $\Delta\nu/\nu=-(30.7\pm 1.7)\times 10^{-19}$.
Excess micromotion – An improved electrical design for our Paul trap significantly reduced the uncompensatable excess micromotion (EMM), which was the largest systematic uncertainty in the previous NIST 27Al+ clock BrewerAccuracy2019 . We optimize micromotion compensation voltages at the start of each day of clock operation, and measure again at the end of each day. These measurements constrain the EMM-induced second-order Doppler shift on the clock transition to $\Delta\nu/\nu=(-1.6\pm 1.6)\times 10^{-19}$; see supplemental material SM for more details.
@@ I Secular motion
We measure the secular mode temperature for all six modes using sideband thermometry, with results shown in Fig. 2b of the main text. For a thermal state, the occupation probabilities $p_{n}$ and the average quantum number $\bar{n}$ are related by $\frac{p_{n+1}}{p_{n}}=\frac{\bar{n}}{\bar{n}+1}$. For red and blue sideband transition probabilities $P_{r}$ and $P_{b}$, this implies $\bar{n}=\frac{P_{r}}{P_{b}-P_{r}}$ LeibfriedQuantumDynamics2003 ; WinelandExperimentalIssues1998 ; TurchetteHeating2000 . As long as the ion’s motion is in a thermal state, this method is valid under a broad range of experimental conditions. One challenge is that the difference $P_{b}-P_{r}$ can be small for $\bar{n}>1$, leading to large statistical fluctuations. We counteract this by taking sufficiently large data sets and by repeating the measurement over the course of several months to check for reproducibility. Fig. S5 shows mode temperature data for all six modes.
@@
To resolve directional ambiguity, we measure $\langle B_{AC}^{2}\rangle$ with the magnetic field and cooling oriented in three different conditions. We label these as $Q$ (original quantization axis, data in first panel of main text Fig. 4A), $H$ (horizontal, second panel), and V ($\sim 10\degree$ from vertical, third panel). The quantization axes $Q$ and $H$ are labeled in Fig. S6, while the axis $V$ is out of plane. In each of the three quantization field configurations, the hyperfine frequency shift due to the ac magnetic field $\langle B_{AC}^{2}\rangle$ is measured using interleaved spectroscopy of two ${}^{25}\mathrm{Mg}^{+}$ hyperfine transitions via the method presented in BrewerMagconst2019 . For each quantization direction, we repeat this measurement at several trap RF powers, yielding the data shown in Fig. 4A of the main text.
@@ III Cooling light Stark shift
$S_{rad}=\frac{2\Delta g_{B}\sqrt{R_{RB}}}{\Gamma^{2}g_{R}g_{rad}^{2}}\frac{(\Omega_{rad}^{meas})^{2}}{\Omega_{co}^{meas}}.$ (5)

### v2

# High-Stability Single-Ion Clock with $5.5\times 10^{-19}$ Systematic Uncertainty

@@ Abstract
Introduction – Optical atomic clocks based on spectroscopy of dipole-forbidden electronic transitions in isolated, trapped atoms are among the most precise instruments developed, capable of measuring time more precisely than the cesium clocks that currently define the second LudlowOpticalClockReview2015 . Accordingly, optical clock frequency ratios are some of the most accurate measurements BeloyBacon2021 ; HausserIndiumMultiIon2025 , and are used as probes for new physics SafronovaNewPhysicsReview2018 , including time variation of fundamental constants SafronovaFundamentalConstantVariation2019 ; SherrillFundamentalConstantVariation2023 ; violations of local position invariance LangePositionInvariance2021 ; constraints on dark matter BeloyBacon2021 ; KennedyDarkMatterClock2020 ; FilzingerClocksDM2023 ; and general relativity at small scales ChouAlClockRelativity2010 ; BothwelStrontiumEnsembleRelativity2022 . Optical clocks based on single trapped ions BrewerAccuracy2019 ; ZhiqiangLutetiumComparison2023 and neutral atoms in optical lattices AppeliStrontiumAccuracy024 have reached fractional frequency uncertainties below $10^{-18}$; further advances open new possibilities for these investigations. Additionally, as the scientific community moves towards the redefinition of the second, advances in the state of the art for clock accuracy and stability are critical DimarcqSecondRoadmap2024 .
@@
We probe both the $\mathrm{m_{F}}=+5/2$ and $\mathrm{m_{F}}=-5/2$ ${}^{1}\mathrm{S}_{0}\leftrightarrow^{3}\mathrm{P}_{0}$ transitions, and generate a “virtual” first-order magnetic-field insensitive transition from their mean frequency RosenbandAlClockObservation2007 . Additionally, we alternate probing from opposite directions, with the two probe beams counterpropagating through the same single-mode optical fibers; the average of opposite directions is insensitive to possible first-order Doppler shifts due to ion motion. Both probe laser acousto-optic modulators (AOMs) are switched on during each cycle, with the inactive direction detuned by 100 kHz. This maintains a nearly identical electric- and magnetic-field environment for probes in either direction and either side of the 27Al+ Zeeman structure.
Decoherence of the 27Al+ clock laser ($\lambda_{L}\simeq 267$ nm) limited previous generations to a 150 ms probe time. We extend the laser coherence by transferring the stability of the JILA cryogenic silicon cavity MateiSiCavity2016 to the clock laser. An ultrastable laser locked to this cavity travels over a 3.6 km path-length-stabilized fiber link to NIST. This fiber link via the Boulder Research and Administrative Network has been described previously BeloyBacon2021 ; for this work, it has been modified as shown in Fig. 1A, such that the cavity-stabilized light serves as reference for an Er:Yb frequency comb FortierCombReview2019 ; NardelliErYbComb2023 . We first pre-stabilize the 27Al+ clock laser with a Pound-Drever-Hall lock to a room-temperature ultra-low expansion glass cavity. With a lower bandwidth of $\sim$10 kHz, we then modulate the drive frequency of an AOM to stabilize the beatnote between clock laser and frequency comb. With this improved clock laser stabilization, we can operate the clock using a 1 s probe duration without substantial loss of atom-laser coherence.
@@
We measure the 27Al+ clock stability by comparison to the JILA Sr optical lattice clock AeppliJILASrClock2024 , with results shown in Fig. 1B. The optical lattice clock fractional stability is $<1\times 10^{-16}/\sqrt{\tau/\mathrm{s}}$, meaning the comparison stability of $3.5\times 10^{-16}/\sqrt{\tau/\mathrm{s}}$ is dominated by the 27Al+ clock. To our knowledge, this represents the lowest instability of any ion clock reported to date. In the future, this stability could be improved by further extending the probe time using differential spectroscopy KimDiffspec2023 , or by extending the clock to multiple spectroscopy ions CuiScalableQLS2022 ; KellerMultIonConf2024 ; PelzerMultiIonCa2024 .
@@
Secular motion – The relativistic time-dilation shift, or second-order Doppler shift, due to secular motion in the ion trap WinelandLaserCoolingLimits1987 is the largest source of systematic uncertainty in this clock. We evaluate the energy in the secular modes of the two-ion crystal using sideband thermometry LeibfriedQuantumDynamics2003 ; WinelandExperimentalIssues1998 . To quantify the statistical uncertainty and test the repeatability of sideband thermometry at the Doppler limit, we repeat the measurements many times over the course of several months. The results are consistent within the measurement uncertainty with the calculated Doppler limit for our nominal laser geometry and secular mode frequencies. Figure 2 shows the record of measurements for one of the radial “stretch” modes, with the others shown in supplemental material SM .
@@
Blackbody radiation – The 27Al+ clock transition has one of the lowest polarizabilities due to blackbody radiation of any existing atomic clock SafronovaBBRclocks2012 . This was recently measured WeiPolarizability2024 with high precision in a 27Al+-40Ca+ clock, using the Stark shift on the 40Ca+ clock transition as a reference for the intensity of a Stark shifting infrared laser and giving a polarizability of $(6.86\pm 0.23)\times 10^{-42}$ $\mathrm{Jm}^{2}/\mathrm{V}^{2}$. We measure the temperature of our apparatus using a set of six thermocouples - three located on the trap and support structure, and three on the surrounding vacuum chamber - constraining the ion’s blackbody environment to be $24.0\pm 3.3$$\degree$°C, for a Stark shift due to blackbody radiation of $\Delta\nu/\nu=-(30.7\pm 1.7)\times 10^{-19}$.
Excess micromotion – Excess micromotion (EMM) was the largest systematic uncertainty in the previous NIST 27Al+ clock BrewerAccuracy2019 . This is substantially reduced due to an improved Paul trap electrical design, featuring a thicker diamond wafer, thicker sputtered gold traces, and rerouting of the traces to balance capacitances on opposite RF electrodes. We optimize micromotion compensation voltages at the start of each day of clock operation, and measure again at the end of each day. These measurements constrain the EMM-induced second-order Doppler shift on the clock transition to $\Delta\nu/\nu=(-1.6\pm 1.6)\times 10^{-19}$; see supplemental material SM for more details on trap design and EMM constraints.
@@
Data are available from the authors upon reasonable request.
@@ I Secular motion
We measure the secular mode temperature for all six modes using sideband thermometry, with results in Fig. 2b above. For a thermal state, the occupation probabilities $p_{n}$ and the average quantum number $\bar{n}$ are related by $\frac{p_{n+1}}{p_{n}}=\frac{\bar{n}}{\bar{n}+1}$. For red and blue sideband transition probabilities $P_{r}$ and $P_{b}$, this implies $\bar{n}=\frac{P_{r}}{P_{b}-P_{r}}$ LeibfriedQuantumDynamics2003 ; WinelandExperimentalIssues1998 ; TurchetteHeating2000 . As long as the ion’s motion is in a thermal state, this method is valid under a broad range of experimental conditions. One challenge is that the difference $P_{b}-P_{r}$ can be small for $\bar{n}>1$, leading to large statistical fluctuations. We counteract this by taking sufficiently large data sets and by repeating the measurement over the course of several months to check for reproducibility. Fig. S5 shows mode temperature data for all six modes.
@@
To resolve directional ambiguity, we measure $\langle B_{AC}^{2}\rangle$ with the magnetic field and cooling oriented in three different conditions. We label these as $Q$ (original quantization axis, data in first panel of main text Fig. 4A), $H$ (horizontal, second panel), and V ($\sim 10\degree$ from vertical, third panel). The quantization axes $Q$ and $H$ are labeled in Fig. S6, while the axis $V$ is out of plane. In each of the three quantization field configurations, the hyperfine frequency shift due to the ac magnetic field $\langle B_{AC}^{2}\rangle$ is measured using interleaved spectroscopy of two ${}^{25}\mathrm{Mg}^{+}$ hyperfine transitions via the method presented in BrewerMagconst2019 . For each quantization direction, we repeat this measurement at several trap RF powers, yielding the data shown in Fig. 4A.
@@ III Cooling light Stark shift
$S_{rad}=\frac{4\Delta g_{B}\sqrt{R_{RB}}}{\Gamma^{2}g_{R}g_{rad}^{2}}\frac{(\Omega_{rad}^{meas})^{2}}{\Omega_{co}^{meas}}.$ (5)